n a certain​ country, the true probability of a baby being a girl is 0.467. Among the next four randomly selected births in the​ country, what is the probability that at least one of them is a boy​?




1
Expert's answer
2021-11-12T08:14:32-0500

Probability that among the next four randomly selected births in the country baby being girl is "0.467^4=0.047562812"

In another way will born at least 1 boy, it means among the next four randomly selected birth in the country probability that at least one of them is a boy "= 1-(0.467)^4"

=1-0.047562812

= 0.952437188 ≈ 0.952